Calculate the amplification factor of a triode valve which has plate resistance of 2 kΩ and transconductance of 2 millimho.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. r base p = 2 KΩ = 2 * 10^3 Ω g base m = 2 milli mho = 2 * 10^–3 mho μ = r base p * g base m = 2 * 10^3 * 2 * 10^–3 = 4 Amplification factor is 4.